Grade 12th passMechanics

hello, I have a question regarding kinematics.
A block of mass 5 Kg is moving on a smooth surface with 20 m/s and ends at a quarter verticle circular loop of radius 2 m. What is the time taken by the block to complete the circular track?

Abhijeet Peshkar
10 Years agoGrade 12th pass

1 Answer

Vikas TU
10 Years ago
taking U = 0 as refernce from smooth surface (horizontal)
conserve K.E and P.E for block.
Therefore,
0.5*5*400 = 5*g*2 + 0.5*5*v^2
v = root(360)
v = 18.973 m/s
 
From 2nd law of eqn.,
2 = 20*t – 0.5g*t^2
1 = 10t –  2.5t^2
or
2.5t^2 – 10t + 1 = 0
solve the quaratic and get the value if t.